OpenStudy (anonymous):

Which of the following is the inequality 2x + 4y < 8 when solved for y? Answer y < 8 - 2x y > 8 -2x y < 2 - 1/2 x Y < 4 - 1/2x

OpenStudy (ja1):

Ok so you need to isolate Y

OpenStudy (anonymous):

how do i do that?

OpenStudy (ja1):

\[2x + 4y < 8 \]

OpenStudy (ja1):

now you subtract -2x from both sides and you get: \[4y < 8 - 2x\]

OpenStudy (anonymous):

why -2x?

OpenStudy (ja1):

Well because the original equation had -2x first and you want to solve for Y not X so you move x to the other side

OpenStudy (ja1):

Now you divide 4 on both sides: \[y < 2 -2x\]

OpenStudy (ja1):

you with me so far?

OpenStudy (anonymous):

yea kindda, after you finish im kindda gonna sit here and understand it all lol

OpenStudy (ja1):

xD well basically you need to get Y alone so you take away the 4 anyways moving on...

OpenStudy (jhannybean):

@charlieberzak It all depends on which variable (either X or Y) you're solving for, to further expand on @JA1's explanation of why you're subtracting -2x

OpenStudy (ja1):

My bad you actually divide 4 on al problems like so: 4y<8−2x /4 /4 /4 y < 2 - 1/2x

OpenStudy (ja1):

So now you get \[y < 2 - 1/2x\] Thanks @jhannybean :)

OpenStudy (anonymous):

thx ja and jhanny :) now its time to sit and stare

OpenStudy (ja1):

So now just sit there and make sense of it all xD

OpenStudy (jhannybean):

\[2x+4y<8\]\[-2x +2x+4y < 8-2x\]\[4y < 8-2x\]\[\frac{4y}{4}< \frac{8-2x}{4}\]\[y< \frac{8}{4}-\frac{2x}{4}\]\[y< 2- \frac{x}{2}\]

OpenStudy (anonymous):

haha exactly

OpenStudy (ja1):

Good luck staring :)

OpenStudy (ja1):

Nice @Jhannybean

OpenStudy (jhannybean):

NO i just put your explanation all together :)

OpenStudy (anonymous):

yep :)

OpenStudy (ja1):

yah but yours was organized xD

OpenStudy (jhannybean):

Pff.

OpenStudy (ja1):

I'm so dead right now I don't know how I pulled it off xD

OpenStudy (anonymous):

lol

OpenStudy (jhannybean):

:) cause youre smart.